OpenStudy (anonymous):

9a-b-2a-10b a) -7a+11b b) 11a+9b c) -1a-9b d) 7a-11b i was thinking a ? i really need help with this

OpenStudy (prettygirl_shynice):

okay ill help u

OpenStudy (anonymous):

okay thank you

OpenStudy (prettygirl_shynice):

ill explain

OpenStudy (shadowlegendx):

Looks like all youʻre doing is like terms

OpenStudy (prettygirl_shynice):

frisr Let's simplify step-by-step. 9a−b−2a−10b =9a+−b+−2a+−10b sencond Combine Like Terms: =9a+−b+−2a+−10b =(9a+−2a)+(−b+−10b) =7a+−11b third answer is==7a−11b

OpenStudy (shadowlegendx):

9a-b-2a-10b Group like terms 9a - 2a - 10b - b 9a - 2a = 7a -10b - b = -11b So we have 7a -11b

OpenStudy (shadowlegendx):

So no, it isnʻt a

OpenStudy (prettygirl_shynice):

yea ur finnal answer 7a-11b and @ShadowLegendX yea

OpenStudy (anonymous):

okay i think i get it better now thank you so much i would have got it wrong

jimthompson5910 (jim_thompson5910):

You can think of it like this (9 apples) - (2 apples) = 9a - 2a = 7a = (7 apples) so if we start off with 9 apples and take away 2 apples, we're left with 7 apples

OpenStudy (prettygirl_shynice):

okay so its d or a

OpenStudy (anonymous):

d right?

OpenStudy (prettygirl_shynice):

u can choose either one okay

OpenStudy (shadowlegendx):

Apples... Iʻm hungry >.>

OpenStudy (anonymous):

lol

OpenStudy (prettygirl_shynice):

lol okay if i were u i would sya d

OpenStudy (anonymous):

thannk you

OpenStudy (anonymous):

gtg

OpenStudy (prettygirl_shynice):

and goodjob